OpenStudy (anonymous):

Compare and contrast simplifying the numeric expression 3 – (2 + 5) to the algebraic expression 3x – (2x + 5)

hero (hero):

So first, just simply solve both and then it become obvious. The expression 3-(2+5), you can evaluate what's inside the parentheses first and then your answer will be in the form of a integer: 3-(2+5) = 3-7 = -4 With the second one, your answer will be in the form of a simplified expression because we don't know what x is: 3x-(2x+5) = 3x-2x-5 = x - 5

hero (hero):

Also, in the second one, you have to expand the parentheses in order to simplify it.